OpenStudy (anonymous):

Find the x-intercept of 3x + 2y = 12 is (__, 0).

OpenStudy (hoblos):

replace y=0

OpenStudy (hoblos):

x=4

OpenStudy (ash2326):

equation of line with x intercept a and y intercept b x/a+y/b=1 here we have 3x+2y=12 3x/12+2y/12=1 x/4+y/6=1 so x intercept 6

OpenStudy (hoblos):

@ash2326 6 is the y-intercept